OpenStudy (anonymous):

A chart with 2 points (-8,-10) and (6,-4). What is the length between the two points?

OpenStudy (anonymous):

\[\sqrt{(6-(-8))^2+(-4-(-10))^2}\] \[=\sqrt{196+36}=\sqrt{232}\]

OpenStudy (anonymous):

okay I am lost as to how you got that as the answer, is there anyway that you can explain it a little bit?

OpenStudy (anonymous):

The formula for distance between two pts (x,y) and (a,b) is as under:\[d=\sqrt((x-a)^2+(y-b)^2)\]

OpenStudy (anonymous):

okay, that helps tremendously thank you